A $10000 investment returns this irregular cash flow as shown in the table. The periodic interest...

  1. A $10000 investment returns this irregular cash flow as shown in the table. The periodic interest rate is 3%.

Dates

15-Jan-2018

16-Feb-2018

5-Mar-2018

14-Jun-2018

11-Dec-2018

Cash Flow

-10000

2000

2500

3500

3500

Calculate the present and future value of this cash flow.
Answer #1

1 investment date = 15-jan-2018, we assume that this is our today's date and pv of initial investment is 10000

we will receive our first cash flow on 16-feb-2018 so period will be 32 days

Present value = Cashflow/ {1+(rate/360)}n   there are 360 days in year n is number of days

=2000/ {1+(0.03/360}32

=2000/ 1.002670

=1994.67

  

we will receive our second cash flow on 5-march-2018 so period will be 49 days

Present value = Cashflow/ {1+(rate/360)}n

=2500/ {1+(0.03/360}49

=2489.81

we will receive our third cash flow on 14-jun-2018 so period will be 150 days

Present value = Cashflow/ {1+(rate/360)}n

= 3500/ {1+(0.03/360}150

=3500/1.012577

=3456.524

we will receive our third cash flow on 11-dec-2018 so period will be330 days

Present value = Cashflow/ {1+(rate/360)}n

= 3500/ {1+(0.03/360}330

=3405.056

FUTURE VALUE our future date will be last cash flow date that is 11 dec 2018

we will receive our first cash flow on 16-feb-2018 so future period will be 298 days

Future  value  = Cashflow* {1+(rate/360)}n

=2000* {1+(0.03/360}298

  = 2000*1.025143

=2050.28

.

we will receive our second cash flow on 5-march-2018 so future period will be 281 days

Future  value  = Cashflow* {1+(rate/360)}n

   =2500* {1+(0.03/360}281

= 2500*1.02369

= 2559.22

we will receive our third cash flow on 14-jun-2018 so future period will be 180 days.


Future  value  = Cashflow* {1+(rate/360)}n

  =3500* {1+(0.03/360}180

=3500*1.01511

=3552.89

future value of forth cash flow is 3500.

Future value of initial investment  = Cashflow* {1+(rate/360)n

=10000* {1+(0.03/360}330

=10000*1.02788

=10278.8